# The dynamical and thermodynamic effects of turbulence for the cosmic baryonic fluid

### v1

@@ 3 Results
We use IllustrisTNG50-1 (hereafter TNG50) simulation data (Pillepich et al., 2018; Nelson et al., 2019a, b; Pillepich et al., 2019) for our work, which is well-suited for studying cosmic fluid turbulence (Wang & He, 2024a, b). All the technical details can be found in Section 2. Generally, turbulence can exert both dynamical and thermodynamic influences on the cosmic fluid, and we will first discuss the dynamical effects.
@@
In Figure 1, we present the mean divergence $\overline{d({\bf x})}$ of the baryonic velocity field as a function of dark matter density $\Delta_{\rm dm}$. We observe that $\overline{d}>0$ in the low-density range, indicating that the fluid is statistically diverging or expanding, i.e., there is a net outflow of fluid from the spatial region. Conversely, $\overline{d}<0$ in the intermediate- and high-density ranges, indicating that the fluid is statistically converging or being compressed, i.e., there is a net inflow of fluid into the region. In the extremely high-density range, which corresponds to small-scale structures such as galaxies, for the highest redshift $z=4$, there is a rapid decrease in divergence, which may indicate an inflow due to rapid star formation activities. For lower redshifts, the divergences undergo rapid growth, which can be attributed to outflows driven by strong astrophysical processes, such as supernova (SN) feedback or active galactic nucleus (AGN) activities (Magorrian et al., 1998; Silk & Rees, 1998; King & Pounds, 2015; Harrison et al., 2018; Sharma & Nath, 2013; Nelson et al., 2019c; Clavijo-Bohórquez et al., 2024). These outflows increase from $z=2$ to $z=1$, nearly remain constant down to $z=0.5$, and then decrease until the present.
@@
In Figure 2, we present the mean dynamical effects as a function of $\Delta_{\rm dm}$. These dynamical $Q$-terms are computed using equations (1) and (2.1). The physical effects of these $Q$-terms are evident: when $Q>0$, the corresponding effect drives the spatial region or cosmic structure to expand, whereas when $Q<0$, it drives the region/structure to collapse. We observe that the cosmic expansion effect $\overline{Q}_{\rm exp}$ simply dampens the flow of the fluid – it suppresses the outflow of gas in both low- and extremely high-density regions (except for the highest redshift $z=4$), and suppresses the inflow of gas in intermediate- and high-density regions. For all redshifts, the gravitational effect $\overline{Q}_{\rm grav}$ tends to expand the gas in low-density regions while compressing the gas in all other regions. Generally, except in high-density regions at $z=0$, $\overline{Q}_{\rm exp}$ and $\overline{Q}_{\rm grav}$ are secondary dynamical effects compared to the turbulent effect $\overline{Q}_{\rm turb}$ and the thermal effect $\overline{Q}_{\rm th}$.
@@
Below, we will specifically analyze the turbulent effect $\overline{Q}_{\rm turb}$, the thermal effect $\overline{Q}_{\rm th}$ and the total combined effect $\overline{Q}_{\rm tot}$ of the gas. We observe that for all redshifts, $\overline{Q}_{\rm turb}<0$ in all regions, and $\overline{Q}_{\rm th}<0$ in low-density regions, but $\overline{Q}_{\rm th}>0$ in all other regions. In both low- and intermediate-density regions, $\overline{Q}_{\rm tot}<0$ and is dominated by $\overline{Q}_{\rm turb}$. The dynamical effects prevent the outflow of the gas in low-density regions and promote the inflow of the gas in intermediate-density regions, causing the gas to be detained in these regions. As previously addressed in Wang & He (2024b), in high-density regions, the dynamical effects of turbulence and thermal pressure are exactly opposite: $\overline{Q}_{\rm th}>0$ while $\overline{Q}_{\rm turb}<0$. However, the total combined effect $\overline{Q}_{\rm tot}>0$ aligns with the thermal effect $\overline{Q}_{\rm th}$, thereby suppressing the inflow of the gas. We cannot account for the outflow in the extremely high-density regions, as we have not included AGN or SN feedbacks into the dynamical equation (1).
@@
Notice that $\overline{Q}_{\rm turb}$ actually represents the combined effects of both bulk and turbulent flow. As analyzed in Wang & He (2024b), bulk flow dominates over turbulent flow at high redshifts, while turbulence becomes increasingly dominant as redshift decreases. At $z=0$, the bulk flow exclusively affects low-density regions, whereas turbulence prevails in all other regions.
@@
Figure 4: Cumulative mass fraction as a function of dark matter density. The cumulative mass fraction $F_{\rm x}$ for hot gas (a), WHIM (b), warm gas (c) and cold component (d) are presented, respectively. The total cumulative mass fraction $F_{\rm tot}$ is shown in Panel (e). Both $F_{\rm x}$ and $F_{\rm tot}$ are functions of dark matter density $\Delta_{\rm dm}$ and redshift $z$. The subscript ‘x’ can be replaced with ‘hot’, ‘WHIM’, ‘warm’ or ‘cold’ as required.
Figure 5: Scatter plot of baryon fraction as a function of dark matter density. Panels (a) to (f) are for $z=4$ to 00, respectively. The horizontal dashed line indicates the cosmic mean value of baryon fraction, $\overline{f}_{\rm b}=0.1573$.
Refer to caption Refer to caption Refer to caption
Figure 6: Mean energy densities and energy ratios as a function of dark matter density. Panels (a) and (b) show the mean energy densities of thermal and turbulent energy at different redshifts, respectively. Panel (c) shows the energy ratios of turbulence to thermal energy. The lines represent the mean values, and the color bands indicate the standard deviations of the values above and below the mean, which are calculated separately. For clarity, only the deviations for $z=0$ are shown; the results for other redshifts are analogous to those of $z=0$.
In Figure 3, we illustrate how the mass fractions of the 4 baryonic components evolve as a function of redshift. We observe that both the cold component and hot gas gradually increase from near zero at $z=4$ to 8.8% and 4.3%, respectively, at $z=0$. However, the most notable change is the significant decrease in the warm gas fraction from 96% to 44%, while the WHIM fraction increases from 3.3% to 43% at the present time. This suggests that there must be some heating mechanism responsible for such a transition. We will address this issue using the cumulative mass fraction, a function of dark matter density $\Delta_{\rm dm}$, which is employed to measure the mass distribution within regions or structures where the density exceeds $\Delta_{\rm dm}$.
Figure 4 displays the cumulative mass fraction $F_{\rm x}$ for the 4 baryonic components as well as $F_{\rm tot}$ for the total matter as a function of the reverse $\Delta_{\rm dm}$. It is observed that the $z$-evolution of these $F_{\rm x}$ is consistent with that shown in Figure 3, and we focus solely on the case at $z=0$. We note that $F_{\rm hot}$ quickly rises to 3.7% at $\Delta_{\rm dm}=200$ and then gradually increases to the saturation fraction of 4.3% towards the lowest-density regions. This suggests that hot gas is predominantly concentrated in virialized structures and slightly extends towards low-density regions. For the cold component, $F_{\rm cold}$ swiftly reaches 8.6% at $\Delta_{\rm dm}=10^{3}$ and remains nearly constant thereafter, indicating that the cold component is entirely confined within galaxies. Regarding the WHIM and warm gas, both $F_{\rm WHIM}$ and $F_{\rm warm}$ increase gradually with decreasing dark matter density, extending to the lowest-density regions with an increasingly gentle growth rate. However, the spatial distribution of the warm gas is more extended than that of the WHIM. Examining the $z$-evolution of $F_{\rm tot}$, we find that for $\Delta_{\rm dm}>200$, baryonic matter continues to accumulate with time, reaching 28% at $\Delta_{\rm dm}=200$ by $z=0$. This suggests that over 70% of baryons at the present time reside in outer regions where $\Delta_{\rm dm}<200$.
In Figure 5, we present the baryon fraction for the total gas of different redshifts. At $z=4$, we see that most of the baryons reside in intermediate-density regions, with baryon fractions centered around the cosmic mean value of $\overline{f}_{\rm b}$. Consistent with the above analysis, as time progresses, dark matter rapidly concentrates and may even form virialized structures, but baryonic matter increasingly decouples from dark matter, leading to a more extended spatial distribution (He et al., 2004). Consequently, baryon fractions are relatively low in high- and extremely high-density regions but are high in low- and intermediate-density regions, as seen in the panels of low redshifts. This accounts for the presence of the ‘high baryon fraction phase’ gas (He et al., 2005).
Figure 7: $z$-evolution of environment-dependent wavelet energy spectra of turbulence. Panels (a), (b) and (c) show the energy spectra of three intervals of dark matter density $\Delta_{\rm dm}$ as indicated in the figure. The energy spectra $E_{\rm turb}(k,\delta)$ are obtained from the env-WPS in equation (15) as $E_{\rm turb}(k,\delta)=k^{2}S_{\rm turb}(k,\delta)$. The relevant power-law fits for the scale ranges that are smaller and larger than the peak positions $k_{\rm peak}$ are indicated in the figure.
We now examine the thermodynamic mechanisms underlying the behaviors of baryonic gas. In this study, we calculate the energy densities of turbulence and thermal energy at different redshifts and analyze their $z$-evolution. This approach allows us to infer the heating and radiative cooling rates of the baryonic fluid. Figure 6 displays the mean thermal energy densities $\overline{\varepsilon}_{\rm th}$, turbulent energy densities $\overline{\varepsilon}_{\rm turb}$, and their mean ratios $\overline{r}$ of the baryonic fluid at different redshifts as a function of $\Delta_{\rm dm}$. As illustrated in Figures 6(a) and (b), both $\overline{\varepsilon}_{\rm th}$ and $\overline{\varepsilon}_{\rm turb}$ increase with $\Delta_{\rm dm}$, yet their evolution from $z=1$ to $z=0$ is extremely weak, indicating that the cosmic fluid is essentially in a state of fully-developed turbulence in this redshift range. The measurements in Figure 6(c) reveal that at low redshifts of $z\leq 1$, the mean ratios $\bar{r}$ are approximately twice as high in both low- and extremely high-density regions. Conversely, in intermediate- and high-density regions, the ratios fall below one, with a minimum value of one-half observed at $\Delta_{\rm dm}\sim 200$ for $z=0$. The $z$-evolution of the ratios is not significant at low redshifts, as the differences in ratios range only between 1 to 2 times between $z=1$ and $z=0$.
At $z=0$, if the dissipation of turbulent and thermal energy is roughly in balance, it is reasonable to expect that the ratio of energy densities, $\overline{r}$, would be approximately 1. However, as aforementioned, there is a deviation from 1, and the reasons for this deviation are as follows. In low- and extremely high-density regions, $\overline{d}>0$, and the fluid can be approximated as adiabatically expanding, which leads to a decrease in temperature $T$ and a drop in thermal energy density $\varepsilon_{\rm th}$. Accordingly, we know from equations (9) and (12) that $\overline{r}$ will increase. Conversely, in intermediate- and high-density regions, $\overline{d}<0$, then the fluid can be approximated as adiabatically compressing, which leads to an increase in $T$ and an increase in thermal energy density $\varepsilon_{\rm th}$, thereby decreasing $\overline{r}$.
@@
Below, we examine the turbulent energy spectrum of the cosmic fluid. Figure 7 displays the $z$-evolution of the environment-dependent wavelet energy spectrum of turbulence. We observe that, while these spectra are very similar to each other, there are indeed some differences. Specifically, the amplitudes of the spectra increase with higher environmental density and generally grow as redshift decreases. However, it should be noted that from $z=1$ to $z=0$, the shapes and amplitudes of the spectra remain essentially unchanged. This suggests that the injection and dissipation of turbulent energy have essentially reached a balance, indicating once again that the fluid is in a state of fully-developed turbulence within this redshift range. Additionally, within the inertial range where $k>k_{\rm peak}$, the power index becomes increasingly flat as redshift decreases and environmental density increases. The power index is $-1.62$ for the case of $\Delta_{\rm dm}\in[8,\infty)$ at $z=0$ (see Figure 7(c)), which is close to the Kolmogorov index of $-5/3$. As analyzed in Wang & He (2024a), the energy dissipation rate for turbulence with a steeper power index than $-5/3$ is higher than that of Kolmogorov turbulence.
@@
We show the scatter plot of gas temperature versus $\Delta_{\rm dm}$ in Figure 8. At $z=4$, we observe that the majority of baryons are in the form of warm gas. As redshift decreases, an increasing amount of warm gas is heated and transitions into the WHIM, with some even transforming into hot gas. Consequently, the mass fraction of warm gas gradually decreases, while that of the WHIM gradually increases. In the case of $z=0$, we find that the WHIM is distributed across nearly all regions but is predominantly concentrated in intermediate-density regions. These findings are consistent with the results shown in Figure 4 and Figure 3.
We discuss two previous works that are relevant to the thermodynamic effects of turbulence in cosmic fluids. The first work, utilizing the WIGEON cosmological hydrodynamic simulations (Feng et al., 2004), Zhang et al. (2006) shows that the cosmic baryonic fluid in the nonlinear regime behaves like Burgers turbulence, which plays a significant role in converting kinetic energy into thermal energy, thereby heating the gas. Moreover, it appears that heating by Burgers turbulence alone is sufficient to account for the basic features of X-ray emission from baryonic gas in the universe. The second work, through an investigation of the turbulent heating rate and radiative cooling rate in the intra-cluster medium (ICM) of the Perseus and Virgo clusters, Zhuravleva et al. (2014) suggests that turbulent heating is sufficient to offset radiative cooling at each radius of the clusters, implying that turbulence is a powerful thermal source to heat the ICM. These findings support the conclusions of our current work.
@@
Figure 8: Scatter plot of baryonic temperature as a function of dark matter density. Panels (a) - (f) are for $z=4$ to 00, respectively. Gas with $T<10^{5}$ is the warm gas, gas with $T>10^{7}$ K is the hot gas, and gas with temperature between these two ranges is the WHIM.

### v3

@@ 3 Results
In this work we used the TNG50-1 simulation sample, which corresponds to a simulated universe of the size $35\,{h^{-1}{\rm Mpc}}$, from the TNG project (Pillepich et al., 2018b; Springel et al., 2018; Marinacci et al., 2018; Nelson et al., 2018; Naiman et al., 2018; Nelson et al., 2019a). TNG uses the moving-mesh code AREPO, which solves the cosmological gravity-magnetohydrodynamic equations on a dynamically unstructured mesh using a second-order accurate Godunov-type scheme (Springel, 2010), and Bauer and Springel (2012) demonstrates its superiority in accurately modeling turbulence. In addition to gravitational and hydrodynamic calculations, the TNG simulations include a comprehensive set of physical processes, such as stellar formation and evolution, the associated metal enrichment, gas cooling, feedback from stellar wind, SNe, and AGNs, and the magnetic field in cosmic structures (Pillepich et al., 2018a; Nelson et al., 2019a). Given the abundance of the physical processes included in the TNG simulations and AREPO’ʼs excellent performance in hydrodynamical computations, the TNG data are well suited for studying cosmic fluid turbulence (Wang and He, 2024b, c).
We used the fourth-order accuracy piecewise cubic spline scheme (Sefusatti et al., 2016) to assign all the simulation particles into a regular cubic $N^{3}_{g}$ grid, with $N_{g}=1536$, for both dark matter and baryonic matter. Hence, the spatial resolution is $22.8h^{-1}{\rm kpc}$, and the mass resolution for dark matter and baryons is $8.4\times 10^{5}h^{-1}{{\rm M}_{\odot}}$ and $1.6\times 10^{5}h^{-1}{{\rm M}_{\odot}}$, respectively. There are no further subdivisions of space and mass.
In all the following relevant computations, we divide the dark matter density ranges from the minimum of $0.75\times 10^{-4}$ to the maximum of $1.45\times 10^{4}$ into 24 logarithmically spaced bins. The computation involves: (1) identifying all grid cells in the density field that fall within a given bin (i.e., greater than or equal to the lower bound and less than the upper bound), (2) locating the corresponding cells in the physical field, and (3) averaging over those cells.
@@
All the technical details can be found in Sect. 2. Generally, turbulence can exert both dynamical and thermodynamic influences on the cosmic fluid, and we will first discuss the dynamical effects.
@@
In Fig. 1 we present the mean divergence $\overline{d({\bf x})}$ of the baryonic velocity field as a function of dark matter density $\Delta_{\rm dm}$. We observe that $\overline{d}>0$ in the low-density range, indicating that the fluid is statistically diverging or expanding. Conversely, $\overline{d}<0$ in the intermediate- and high-density ranges, indicating that the fluid is statistically converging or being compressed. In the extremely high-density range, which corresponds to small-scale structures such as galaxies, for the highest redshift $z=4$, there is a rapid decrease in divergence, which may indicate gas convergence or compression due to rapid star formation activities. For lower redshifts, the divergences undergo rapid growth, which can be attributed to expansion driven by strong astrophysical processes, such as star formation, SN feedback, or AGN activities (Magorrian et al., 1998; Silk and Rees, 1998; King and Pounds, 2015; Harrison et al., 2018; Sharma and Nath, 2013; Nelson et al., 2019b; Clavijo-Bohórquez et al., 2024). These expansions increase from $z=2$ to $z=1$, nearly remain constant down to $z=0.5$, and then decrease until the present.
Figure 5: Panel (a): Averaged dynamical effects of $z=0$ contributed by turbulence and shocks. Panel (b): Averaged thermodynamic effects of $z=0$ contributed by turbulence and shocks. The shock-finding algorithm of Schaal and Springel (2015) and Schaal et al. (2016) was used to identify shock zones, with a minimum Mach number ($\mathcal{M}_{\rm min}$) of $1.0$.
@@
In Fig. 2 we present the mean dynamical effects as a function of $\Delta_{\rm dm}$. These dynamical $Q$-terms are computed using Eqs. (1) and (2.1). The physical effects of these $Q$-terms are evident: when $Q>0$, the corresponding effect drives the spatial region or cosmic structure to expand, whereas when $Q<0$, it drives the region/structure to collapse. We observe that the cosmic expansion effect $\overline{Q}_{\rm exp}$ simply dampens the flow of the fluid – it suppresses the expansion of gas in both low- and extremely high-density regions (except for the highest redshift, $z=4$), and suppresses the convergence of gas in intermediate- and high-density regions. For all redshifts, the gravitational effect $\overline{Q}_{\rm grav}$ tends to expand the gas in low-density regions while compressing the gas in all other regions. Generally, except in high-density regions at $z=0$, $\overline{Q}_{\rm exp}$ and $\overline{Q}_{\rm grav}$ are secondary dynamical effects compared to the turbulent effect $\overline{Q}_{\rm turb}$ and the thermal effect $\overline{Q}_{\rm th}$.
@@
Below, we specifically analyze the turbulent effect $\overline{Q}_{\rm turb}$, the thermal effect $\overline{Q}_{\rm th}$, and the total combined effect $\overline{Q}_{\rm tot}$ of the gas. We observe that for all redshifts, $\overline{Q}_{\rm turb}<0$ in all regions, and $\overline{Q}_{\rm th}<0$ in low-density regions, but $\overline{Q}_{\rm th}>0$ in all other regions. In both low- and intermediate-density regions, $\overline{Q}_{\rm tot}<0$ and is dominated by $\overline{Q}_{\rm turb}$. The dynamical effects prevent the expansion of the gas in low-density regions and promote the convergence of the gas in intermediate-density regions, causing the gas to be detained in these regions. As previously addressed in Wang and He (2024c), in high-density regions, the dynamical effects of turbulence and thermal pressure are exactly opposite: $\overline{Q}_{\rm th}>0$ while $\overline{Q}_{\rm turb}<0$. However, the total combined effect $\overline{Q}_{\rm tot}>0$ aligns with the thermal effect $\overline{Q}_{\rm th}$, thereby suppressing the convergence of the gas. We cannot account for the expansion in the extremely high-density regions, as we have not included star formation, AGN, or SN feedback in the dynamical equation, Eq. (1).
@@
Note that $\overline{Q}_{\rm turb}$ actually represents the combined effects of bulk and turbulent flow. As analyzed in Wang and He (2024c), bulk flow dominates over turbulent flow at high redshifts, while turbulence becomes increasingly dominant as redshift decreases. At $z=0$, the bulk flow exclusively affects low-density regions, whereas turbulence prevails in all other regions.
We next examined the thermodynamic mechanisms underlying the behaviors of baryonic gas. In this study we calculated the energy densities of turbulence and thermal energy at different redshifts and analyzed their $z$-evolution. This approach enabled us to infer the heating and radiative cooling rates of the baryonic fluid. Figure 3 displays the mean thermal energy densities $\overline{\varepsilon}_{\rm th}$, turbulent energy densities $\overline{\varepsilon}_{\rm turb}$, and their mean ratios $\overline{r}$ of the baryonic fluid at different redshifts as a function of $\Delta_{\rm dm}$. As illustrated in Figs. 3a and 3b, both $\overline{\varepsilon}_{\rm th}$ and $\overline{\varepsilon}_{\rm turb}$ increase with $\Delta_{\rm dm}$, yet their evolution from $z=1$ to $z=0$ is extremely weak, indicating that the cosmic fluid is essentially in a steady state in this redshift range, in which the statistical properties of the cosmic fluid, such as the energy density or energy spectrum, no longer change with time. The measurements in Fig. 3c reveal that at low redshifts of $z\leq 1$, the mean ratios $\bar{r}$ are approximately 2 in both low- and extremely high-density regions. Conversely, in intermediate- and high-density regions, the ratios fall below one, with a minimum value of one-half observed at $\Delta_{\rm dm}\sim 200$ for $z=0$. The $z$-evolution of the ratios is not significant at low redshifts, as the differences in ratios range only between 1 and 2 times between $z=1$ and $z=0$.
If the kinetic energy of the cosmic fluid can be entirely converted into thermal energy through mechanisms such as turbulence or shocks, and there are no other heating mechanisms, or they can be neglected, then this can explain why the ratio $\overline{r}$ roughly equals $1$ at $z=0$. If the dissipation of turbulent energy and thermal energy is roughly in balance, it can explain why the ratio $\overline{r}$ remains essentially constant within the range of 0<z<10<z<1111We note that in Miniati and Beresnyak (2015), this ratio for the ICM is about $1/3$ and also does not change with time. Clearly, while the constancy over time is consistent with our findings, the ratio itself is different from ours. We are not sure of the exact reason for this discrepancy, but we speculate that it may be due to the different numerical methods and baryonic physics processes they have used compared to ours.. However, as aforementioned, the ratio $\overline{r}$ deviates from $1$, and the reasons for this deviation are as follows. In low- and extremely high-density regions, $\overline{d}>0$, and the fluid can be approximated as adiabatically expanding, which leads to a decrease in temperature ($T$) and a drop in thermal energy density ($\varepsilon_{\rm th}$). Accordingly, we know from Eqs. (3) and (6) that $\overline{r}$ will increase. Conversely, in intermediate- and high-density regions, $\overline{d}<0$, then the fluid can be approximated as adiabatically compressing, which leads to an increase in $T$ and an increase in thermal energy density $\varepsilon_{\rm th}$, thereby decreasing $\overline{r}$.
Figure 6: Mass fraction as a function of redshift. The mass fractions of the four baryonic components evolve with redshift.
@@
Figure 7: Cumulative mass fraction as a function of dark matter density. The cumulative mass fraction ($F_{\rm x}$) is presented for hot gas (a), the WHIM (b), warm gas (c), and the cold component (d). The total cumulative mass fraction ($F_{\rm tot})$ is shown in panel (e). Both $F_{\rm x}$ and $F_{\rm tot}$ are functions of dark matter density ($\Delta_{\rm dm}$) and redshift ($z$). The subscript “x” can be replaced with “hot,” “WHIM,” “warm,” or “cold” as required.
@@
Below, we examine the turbulent energy spectrum of the cosmic fluid. Figure 4 displays the $z$-evolution of the environment-dependent wavelet energy spectrum of turbulence. We observe that, while these spectra are very similar to each other, there are indeed some differences. Specifically, the amplitudes of the spectra increase with higher environmental density and generally grow as redshift decreases. However, it should be noted that from $z=1$ to $z=0$, the shapes and amplitudes of the spectra remain essentially unchanged. This suggests that the injection and dissipation of turbulent energy have essentially reached a balance, indicating once again that the fluid is in a steady state within this redshift range. Additionally, within the inertial range where $k>k_{\rm peak}$, the power index becomes increasingly flat as redshift decreases and environmental density increases. The power index is 1.44-1.44 for $\Delta_{\rm dm}>200$ at $z=0$, which is close to the Kolmogorov index of 5/3-5/3. It is 2.41-2.41 for $\Delta_{\rm dm}<0.2$ and 2.25-2.25 for $0.2<\Delta_{\rm dm}<200$. These indices are steeper than that of Kolmogorov turbulence and even steeper than that of Burgers turbulence. We speculate that the steeper indices may be caused by the high inhomogeneity and anisotropy of the cosmic baryonic fluid, which is widely distributed in these regions of relatively low environmental density. Therefore, it does not conform to the results of Kolmogorov and Burgers turbulence, which assume homogeneity and isotropy in spatial distribution.
We used Eq. (2.3) to calculate the dynamical effects of turbulence and shocks at $z=0$, i.e., $\overline{Q}_{\rm turb}$ and $\overline{Q}_{\rm shock}$. From Fig. 5a, we observe that the dynamical effects of both turbulence and shocks are negative, which, as analyzed earlier, indicates that both turbulence and shocks act to compress the gas and hinder its tendency to flow outward. We find that in low-density regions the mean dynamical effects of shocks are marginally smaller than, yet comparable to, those of turbulence. However, in intermediate-density regions, starting from $\Delta_{\rm dm}=0.1$, the dynamical effects of turbulence become progressively stronger with increasing environmental dark matter density. Therefore, we conclude that in intermediate-density regions, the dynamical effects of turbulence dominate over those of shocks. In high-density regions, the dynamical effects of shocks can be completely neglected compared to turbulence.
We used Eq. (2.3) to calculate the thermodynamic effects of turbulence and shocks at $z=0$, i.e., $\overline{\varepsilon}_{\rm turb}$ and $\overline{\varepsilon}_{\rm shock}$. From Fig. 5b, it can be seen that the mean thermal effects of turbulence are much greater than those of shocks. In regions of $\Delta_{\rm dm}<0.5$, the thermal effects of turbulence are roughly 40 times as high as those of shocks, and this ratio further increases with the environmental dark matter density. The ratio exceeds four orders of magnitude when $\Delta_{\rm dm}>200$, and exceeds nearly six orders of magnitude when $\Delta_{\rm dm}>10^{3}$. This implies that in virialized structures, the heating of gas by shocks can be completely neglected compared to that by turbulence. For a long time, shocks have been considered important thermal sources responsible for heating the ICM and the IGM. However, such a perspective needs to be revised. We see that turbulence is a more powerful heating mechanism than shocks. The above conclusions regarding dynamical and thermodynamical effects of the cosmic gas are also similar for other redshifts. If we applied the value of $\mathcal{M}_{\rm min}=1.3$ to identify shock zones, as recommended by Schaal and Springel (2015), then the dynamical and thermodynamic effects of shocks would be even further reduced.
Figure 8: Scatter plot of the baryon fraction of baryonic gas as a function of dark matter density. Panels (a) to (f) are for $z=4$ to 0, respectively. The horizontal dashed line indicates the cosmic mean value of the baryon fraction, $\overline{f}_{\rm b}=0.1573$.
In Fig. 6 we illustrate how the mass fractions of the four baryonic components evolve as a function of redshift. We observe that both the cold component and hot gas gradually increase from near zero at $z=4$ to 8.8% and 4.3%, respectively, at $z=0$. However, the most notable change is the significant decrease in the warm gas fraction from 96% to 44%, while the WHIM fraction increases from 3.3% to 43% at the present time. These evolutionary results should be attributed to the aforementioned heating mechanisms, especially turbulent heating. We delved into this issue using the cumulative mass fraction, a function of dark matter density ($\Delta_{\rm dm}$), which is employed to measure the mass distribution within regions or structures where the density exceeds $\Delta_{\rm dm}$.
Figure 7 displays the cumulative mass fraction $F_{\rm x}$ for the four baryonic components as well as $F_{\rm tot}$ for the total matter as a function of the reverse $\Delta_{\rm dm}$. It is observed that the $z$-evolution of these $F_{\rm x}$ is consistent with that shown in Fig. 6, and we focused solely on the case at $z=0$. We note that $F_{\rm hot}$ quickly rises to 3.7% at $\Delta_{\rm dm}=200$ and then gradually increases to the saturation fraction of 4.3% toward the lowest-density regions. This suggests that hot gas is predominantly concentrated in virialized structures and slightly extends toward low-density regions. For the cold component, $F_{\rm cold}$ swiftly reaches 8.6% at $\Delta_{\rm dm}=10^{3}$ and remains nearly constant thereafter, indicating that the cold component is entirely confined within galaxies. Regarding the WHIM and warm gas, both $F_{\rm WHIM}$ and $F_{\rm warm}$ increase gradually with decreasing dark matter density, extending to the lowest-density regions with an increasingly gentle growth rate. However, the spatial distribution of the warm gas is more extended than that of the WHIM. Examining the $z$-evolution of $F_{\rm tot}$, we find that for $\Delta_{\rm dm}>200$, baryonic matter continues to accumulate with time, reaching 28% at $\Delta_{\rm dm}=200$ by $z=0$. This suggests that over 70% of baryons at the present time reside in outer regions where $\Delta_{\rm dm}<200$.
In Fig. 8 we present the baryon fraction for the total baryonic gas of different redshifts. At $z=4$, we see that most of the baryons reside in intermediate-density regions, with baryon fractions centered around the cosmic mean value of $\overline{f}_{\rm b}$. Consistent with the above analysis, as time progresses, dark matter rapidly concentrates and may even form virialized structures, but baryonic matter increasingly decouples from dark matter, leading to a more extended spatial distribution (He et al., 2004). Consequently, baryon fractions are relatively low in high- and extremely high-density regions but are high in low- and intermediate-density regions, as seen in the panels of low redshifts. This accounts for the presence of the “high baryon fraction phase” gas (He et al., 2005). The evolution of the baryon fraction in such a manner, especially in the intermediate-density regions, is precisely a manifestation of the dynamical suppression effects (negative $\overline{Q}_{\rm turb}$) discussed above, which are primarily contributed by turbulence in the cosmic fluid. For discussions of the relationship between the baryon fraction and dynamical effects, we refer the reader to the earlier analysis of Fig. 3.
@@
We show the scatter plot of gas temperature versus $\Delta_{\rm dm}$ in Fig. 9. At $z=4$, we observe that the majority of baryons are in the form of warm gas. As redshift decreases, an increasing amount of warm gas is heated and transitions into the WHIM, with some even transforming into hot gas. Consequently, the mass fraction of warm gas gradually decreases, while that of the WHIM gradually increases. In the case of $z=0$, we find that the WHIM is distributed across nearly all regions but is predominantly concentrated in intermediate-density regions. These findings are consistent with the results shown in Figs. 6 and 7.
@@
Figure 9: Scatter plot of temperature for baryonic gas as a function of dark matter density. Panels (a) - (f) are for $z=4$ to 0, respectively. Gas with $T<10^{5}$K is the warm gas, gas with $T>10^{7}$ K is the hot gas, and gas with temperatures between these two ranges is the WHIM. Note that regions of temperature below $10^{4}$ K in the TNG data are not well resolved, particularly for the cells of star-forming gas (see Springel and Hernquist, 2003), and should be interpreted with caution.